OpenStudy (anonymous):

Sumbdy tld dat formula for Rth for Q6 is Vth/Isc. Bt hw to calculate Isc???? Can sumbdy help me out pls..????

OpenStudy (anonymous):

1)Vo* (R3-(B*R1*R3)+R1+R2+A*R2)/(R3-B*R1*R3)=VIN 2)i1=V0/(R3-B*R1*R3) 3)vo = vth 4)Rth=Vth/Isc

OpenStudy (anonymous):

But how to calculate Isc ????

OpenStudy (anonymous):

tell me ur values...

OpenStudy (anonymous):

Vin=15V, R1=0.5Ω, R2=1.5Ω, R3=5Ω, A=2 and B=1.5

OpenStudy (anonymous):

Vin = R1 Isc + R2 Isc IsC =Vin/ R1 + R2 Rth = Vth / Isc This is the formula?

OpenStudy (nurali):

Question 6 Formula Vout = Vin*(R3-B*R1*R3)/(R3-(B*R1*R3)+R1+R2+A*R2) i1=V0/(R3-B*R1*R3) and Vth is same as Vout. Isc=vIN*(1-B*R1)/(A*R2+R1+R2) Rth=VO/iSC where VO is Vth

OpenStudy (nurali):

If correct please click best response